Home CMS Production Clean room MedAustron HEPHY testbeams old
testbeam-DESY-Oct2017 testbeam-DESY-April2017 testbeam-SPS2015 testbeam-SPS2014 testbeam-DESY14 testbeam-SPS12 testbeam-SPS11 testbeam-SPS10 testbeam-SPS09 testbeam-SPS08 LP-TPC
  HEPHY logbook for testbeam at SPS June 2009  Not logged in ELOG logo
Message ID: 34     Entry time: Thu Aug 20 18:56:25 2009
Author: Christian Irmler 
Project: alignment_laser 
Subject: laser 
Run Number: laser001 
Events: 100000 
StartTime: 20.08.2009 18:48:53 
EndTime:  

2 alignment modules
location in beam direction: h04-h03
laser is shooting on h03 (opposite to the beam)
 

HV=70V, 40MHz, 6 samples
beam off
module 1 (h03) is used w inverter off
since the signal off the second sensor is only 10 percent, the base line of h03 was increased again (now @ 260 ADC).

5ns trigger pulse width, 1 kHz trigger, 
current_laser=10mA,
Photodiode: pulse_width=15ns, amplitude=7mV

bad strips: APV5: 115, APV6: 1

ELOG V3.1.4-3d7c366